### v1

# Ferrotoroidal order of the monolayer ice in martyite

Abstract
The beauty of snowflakes, the fractal patterns with the hexagonal symmetry, reflects the crystal structure of ice. The honeycomb network of H2O molecules stacks along the perpendicular direction, composing the hexagonal crystalline ice (ice Ih phase). What can one see if one layer of ice is extracted and confined to a two-dimensional (2D) sheet? Martyite Zn3(V2O7)(OH)2$\cdot$2H2O, a mineral with the honeycomb lattice of H2O in the porous framework, is an ideal system to study such monolayer ice. Due to the geometrical frustration and 2D nature of the honeycomb lattice, H2O molecules are dynamically disordered at room temperature, and the most stable molecular arrangement is nontrivial. In this study, we reveal disorder-order transitions of H2O in martyite using the single-crystal x-ray diffraction (XRD) and molecular dynamics (MD) simulation. The XRD results visualize the formation of hydrogen-bonded toroidal H2O hexamers, leading to the electric ferrotoroidal order below 200 K. Combined with the MD simulations, we discuss the formation process of the H2O hexamers and how they compromise the molecular arrangement towards lower temperatures. Our results unveil the ground state of monolayer ice, a fundamental knowledge to understand the polymorphism of H2O.
Main text
Figure 1: Crystal structures of martyite. (a) Crystal structure of martyite at room temperature. Three favoured orientations of H2O are shown for the A and B sublattices of the honeycomb lattice. Dashed lines show the hydrogen bond. (b) AB-site interaction energy scheme for the 3-states model. The stable geometry is arranged from left to right. (c) Ordered structure of H2O molecules in martyite at 30 K. The dotted and solid lines show the HT and LT unit cells, respectively. Representative interatomic distances are also shown. A perspective view of one H2O hexamer is shown in the right panel. (d) Ferrotoroidal order of the monolayer ice in martyite. The blue and green arrows show electric dipole moments of H2O and electric toroidal dipole moments of H2O hexamers, respectively.
The polymorphism of water is a central interest of science. Ice shows diverse forms in different environments, from fluttering snowflakes to dense ice inside the core of planets [1, 2, 3]. The natural ice on the Earth (ice Ih) is a hydrogen-bonded crystal with a stack of buckled honeycomb layers of H2O molecules. The local configuration satisfies the two-covalent and two-hydrogen bonds in the tetrahedral geometry, the so-called ice rules (Bernal–Fowler rules) [4]. Since there are multiple ways to satisfy the ice rules, the proton configurations in ice Ih are disordered with Pauling’s residual entropy [5], reminiscent of the frustrating nature of ice. On the delicate balance of intermolecular interactions, at least 20 phases of ice have been reported in various temperature and pressure ranges [6, 7, 8, 9]. Among them, ices XVI and XVII are porous phases obtained by removing guest molecules from clathrate hydrates [10, 11, 12].
Confinement is another approach to realizing novel states of water [13, 14]. Interfacial water formed on a metal surface or graphene sheets has been studied [15, 16, 17, 18]. Their local structures are similar to ice Ih, satisfying the ice rule, although the periodicity of the substrate globally influences the symmetry. Confined water inside minerals [19, 20, 21, 22, 23], molecular-organic frameworks [24, 25], and carbon nanotube [26, 27] has also been of interest. Here, the intermolecular distance and the coordination geometry can be controlled, extending the diversity of water to be explored.
In this study, we focus on a hydrated mineral, martyite Zn3(V2O7)(OH)2$\cdot$2H2O, with the trigonal crystal system (space group $P\overline{3}m1$) [28, 29]. Martyite has the quasi-2D porous framework, made of the block layers of ZnO4(OH)2 octahedra separated by the V2O7 pillars (Fig. 1(a)). The water of crystallization confined in the porous framework forms the honeycomb network of H2O, which is isostructural to the monolayer ice. Each H2O is coordinated by the hydroxy group of the framework and forms the hydrogen bonds alternately in the ±$\pm c$ directions. This hydrogen bond moderately fixes the oxygen atom of H2O, while, the proton configuration (molecular orientation) can change in the $ab$ plane. As a result, the H2O molecules in martyite are dynamically disordered with the three favoured orientations along the nearest-neighbour directions of the honeycomb network, similar to the plastic ice VII [30].
Our central question is, how these H2O molecules are arranged at low temperatures. As a first approximation, considering the three orientations at the A and B sites of the honeycomb lattice, four types of AB-site configurations are expected (Fig. 1(b)). Here, the configurations are listed in increasing order of energy from left to right, considering the hydrogen bond and dipole-dipole interaction. Note that H2O is a polar molecule with a local electric dipole moment pointing from the O atom toward the midpoint between the two H atoms. We comment that the out-of-plane interaction energy is of the order of 1 % of the in-plane interaction (see supplemental materials, SM [31]). Thus, water in martyite can be regarded as monolayer ice (quasi-2D system), where the AB-site interaction with the hydrogen bond plays a dominant role. On the honeycomb lattice of H2O, the average number of the hydrogen atoms per bonds is 4/3. Namely, all the nearest-neighbour molecules cannot be hydrogen-bonded. Therefore, the orientational degree of freedom of the monolayer ice is geometrically frustrated, leading to a highly degenerate ground state. In this study, we reveal the ground state of monolayer ice and discuss how the system lifts the frustration at low temperatures.
Dielectric properties
Figure 2: Dielectric anomalies of martyite. Temperature dependence of the dielectric constant $\varepsilon^{\prime}$ and dissipation factor $D$ of the polycrystalline martyite.
Figure 2 shows the dielectric constant $\varepsilon^{\prime}$ and dissipation factor $D$ of the polycrystalline martyite as a function of temperature $T$. Two anomalies are observed at $T_{\mathrm{c1}}\approx$ 170–200 K and $T_{\mathrm{c2}}\approx$ 30–50 K. The decreases of $\varepsilon^{\prime}$ indicate the ordering of the H2O molecules rotating at room temperature. Based on the x-ray diffraction (XRD), Raman spectra, and molecular dynamics (MD) simulation, we conclude that these anomalies are related to the two structural phase transitions driven by the disorder-order transition of H2O (see following sections and SM [31]). Accordingly, we term the phases as HT ($T_{\mathrm{c1}}<T$), LT ($T_{\mathrm{c2}}<T<T_{\mathrm{c1}}$), and LT’ ($T<T_{\mathrm{c2}}$). The frequency dependences of $T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$ indicate the dielectric relaxation where the orientational degree of freedom gradually freezes. The observed dynamics reflects the energy barrier for rotational motions of H2O to relax to nearly-degenerate but lower energy states.
Single crystal XRD
Figure 3: Superlattice Bragg reflections of martyite. (a-d) XRD patterns of the single crystal of martyite at (a, b) 300 K and (c, d) 30 K. (e) Temperature dependence of the superlattice Bragg peak intensity at the index of $-4/3$, $5/3$, 0. (f) Temperature dependence of the lattice constants $a$ and $c$ of the original unit cell.
Figures 3(a-d) show the single-crystal XRD patterns of martyite at 300 K and 30 K. We analyzed the XRD data at 300 K and confirmed that the crystal structure is consistent with the previous report [28, 29] (see SM [31]). Between the layers of martyite, the O atoms of H2O molecules form a slightly buckled honeycomb network with equivalent O-O bonds. Each O atom is bonded to three H atoms with a 2/3 occupancy, allowing the H2O molecule to retain the orientation degree of freedom of three-state vector (Fig. 1(a)). Superlattice peaks appear at the wave vector $q=\{1/3,1/3,0\}$ below $T_{\mathrm{c1}}\approx 200$ K, indicating a structural phase transition. Figure 3(e) shows the temperature dependence of the superlattice peak intensity -4/3 5/3 0 as an order parameter. No peak splitting or extinction rules are observed in any of the fundamental and superlattice peaks. Therefore, the unit cell becomes $\sqrt{3}a\times\sqrt{3}a\times c$, and the trigonal symmetry is maintained. We note that no superlattice peaks or diffuse scattering are observed along the $c^{*}$-axis in the low-temperature phase. This indicates that the ordering of H2O molecules is in-phase along the $c$ axis and is long-range ordered in martyite, in contrast to the proton-disordered ice Ih [32].
We conclude that the space group of the LT phase at 30 K is $P\overline{3}$ (for details, see SM [31]). This lowering of symmetry generates merohedral twins due to the disappearance of mirror planes perpendicular to the tertiary axes of the trigonal lattice. In the current crystal, these domains exist in an approximately 1:1 ratio (see SM [31]). Figure 1(c) shows the ordered arrangement of H2O molecules at 30 K. The O atoms of the H2O molecules, which form a honeycomb lattice in the HT phase, are distorted into a breathing honeycomb lattice, forming H2O hexamers in the LT phase. In this hexamer, H2O molecules form a toroidal ring. One H atom in the H2O molecule points toward a neighbouring H2O molecule, while the other H atom is offset from the honeycomb plane. The off-plane H atoms are arranged alternately up and down as coordinated from the ZnO4(OH)2 block layer (Fig. 1(c)). This alternation in the monolayer ice resembles the hydrogen bonds in ice Ih formed in the ±$\pm c$ directions.
Figure 3(f) shows the temperature dependence of the lattice constants. There is no anomaly in the $a$ and $c$ parameters, suggesting that the framework of martyite is rigid against the breathing deformation of the honeycomb network of H2O. The observed superlattice peak intensities are bilinear to the atomic displacements of O of H2O (Fig. S2(f) [31]).
MD simulation
Figure 4: H2O hexamer and octadecamer formation in the MD simulation. (a) The number of hydrogen bonds per water molecule (blue) and the ratio of the type-1 H2O hexamers (red) as a function of temperature. The ratio of the type-1 hexamers is calculated for the designed ferrotoroidal state. (b) The in-plane radial distribution function between O atoms of H2O. The curves are shifted for clarity. (c) The snapshots of the hydrogen bonding network. The arrows represent the hydrogen-bonded pair defined by the intermolecular O–O distance ($d_{\mathrm{OO}}$) and the hydrogen bonding angle ($\theta$) in the right figure. Clockwise (counterclockwise) hexamers are marked by green crosses (yellow circles). (d) Spatial distribution of the hexamers, type 0, 1, and 2, simulated for the designed ferrotoroidal state. As shown in the right figures, purple (0), red (1), and blue (2) hexagons represent the H2O hexamers with zero, one, and two longer hydrogen bonds. The black dots indicate that at least one hydrogen bond in the hexamer is broken. The dotted triangles indicate the H2O octadecamers. The dashed line indicates the domain boundary of the lattice of the octadecamers.
MD simulations on the monolayer ice in martyite demonstrate how the randomly oriented water molecules start to form hydrogen bonds with their neighbours. The simulation details and movie are given in SM [31]. In the simulation, we use the water model TIP4P/ICE [33] and design one layer of martyite, namely, the monlayer ice sandwitched by the ZnO4(OH)2 layers supported by the V2O7 pillars. All the structural parameters are taken from the XRD results. Water molecules interact with each other and with atoms of the framework via Lennard-Jones (LJ) and Coulomb potentials. Although no artificial constraints are imposed on the water molecules, their translational motions are naturally restricted to those around a honeycomb lattice site even at 300 K. No quantum effect on the dynamics of atoms is assumed in the MD simulations.
Figure 4(a) shows the number of hydrogen bonds per water molecule $n_{\rm HB}$ as a function of temperature cooled down from 300 K. The results obtained with the TIP4P/2005 model are presented in SM [31], which are essentially the same as those with the TIP4P/ICE model. At 300 K, $n_{\rm HB}<1$, meaning that each water molecule in martyite assumes an almost random orientation. As the temperature decreases, $n_{\rm HB}$ gradually increases from 220 K, rises sharply between 150 and 200 K and reaches its maximum value of 2 at 100 K. Figure 4(c) depict the representative patterns of hydrogen bonds shown by the arrows at 80 K, 130 K, 180 K, and 300 K. At 300 K, the hydrogen-bonded molecules are sparse, among which the population of dimers is the largest. At the temperature of 180 K, where $n_{\rm HB}$ changes most rapidly, water molecules have one or two hydrogen bonds and together form structures of hexagonal rings and a variety of chains. At lower temperatures, water molecules form a regular array of hydrogen-bonded hexagonal rings. Each hexagonal ring has clockwise (CW) or counterclockwise (CCW) proton ordering as shown by the green crosses and yellow circles, respectively. The formation and breaking of hydrogen bonds takes place at high temperatures, but once the regular array of hexagonal rings is established, the hydrogen bonds are stable and no rearrangement occurs. The simulated transition from the orientationally disordered state (HT) to the array of hexagonal rings (LT) agrees with the XRD results. The transition temperature ($T_{\mathrm{c1}}$) also coincides with the experiment, indicating that the water molecules in martyite are well described by the monolayer ice model.
[…]
Electric ferrotoroidal order
In the H2O hexamer, H2O molecules form the toroidal ring structure. Here, a head-to-tail configuration of the electric dipole moments ($\mathbf{D}$) form an electric toroidal dipole moment defined by $\mathbf{A}=\sum\mathbf{r}\times\mathbf{D}$ (Fig. 1(d)). Our XRD results reveal that the LT phase of martyite shows an electric ferrotoroidal order driven by the disorder-order transition of water. So far, several compounds are known to show the electric ferrotoroidal order (also referred to as ferroaxial or ferro-rotational order, RbFe(MoO4)2, Ca5Ir3O12, and NiTiO3) [34, 35, 36, 37], which are driven by the rotational displacements of the coordination polyhedra. Amongst them, martyite is the simplest case where electric dipoles of polar molecules cooperatively form the toroidal motif, spontaneously breaking the mirror planes parallel to the $c$ axis. Here, the intermolecular interactions are relatively simple, and the MD simulation can reproduce the toroidal ordering. In the simulation, the signs of the toroidal moments (±$\pm\mathbf{A}$) are more or less random since the MD simulation is not long enough at each temperature (10 ns), and therefore the cooling rate is too fast to reach fully equilibrated states. However, once such an initial configuration is given by design, the ferrotoroidal order remains at least up to 100 K, and the internal energy is consistently lower than that of the simulated system without ferrotoroidal order (see SM [31]). The stability of the ferrotoroidal order can also be explained from the AB-site interaction energy scheme (Fig. 1(b) and see also Fig. S6 [31]). Thus, the ground state of the monolayer ice is a ferrotoroidal order, although the simulation gives the random toroidal pattern as a metastable state.
To reveal the fine structure of the ground state, we design the forced ferrotoroidal state (all hexamers are CW) and see how the molecular arrangements change towards lower temperatures. Figures 4(d) show the snapshots of the water hexamers cooled down from the ferrotoroidal state. Here, type 0, 1, and 2 hexagons represent the H2O hexamers with zero, one, and two longer hydrogen bonds ($d_{\mathrm{OO}}>3.35$). At 120 K, various types of hexamers, including the broken ones shown by the black dots, are distributed randomly. By decreasing the temperature, the ratio of type-1 hexamers rapidly increases at $T_{\mathrm{c2}}\approx 100$ K (Fig. 4(a)). Here, three type-1 hexamers form a H2O octadecamer (dotted triangles in Fig. 4(d)). At the lowest temperature (LT’), H2O octadecamers alternatively align along one direction, leading to the orthorhombic distortion from the LT phase.
The hexamers’ deformations are likely caused by the lattice constants of the honeycomb lattice, which are determined by the framework size. Specifically, the distance of 3.36 Å is significantly larger than 2.76 Å, the ideal hydrogen bond length in ice Ih. Thus, the H2O hexamer deforms for the sake of five stronger hydrogen bonds at the expense of one weaker one. Such structural instability is detected by the dielectric and Raman-scattering experiments at around $T_{\mathrm{c2}}\approx$ 30–50 K (Fig. 2 and Fig. S9 [31]).
The MD simulation suggests that the LT’ phase is partially disordered (see the domain boundary shown by the dashed line in Fig. 4(d)). The energy cost to create a domain boundary is almost zero within our calculation accuracy. Besides, the distribution of the deformed hexamers depends on the water model and system size (see SM [31]). Therefore, the ground state of the martyite is highly degenerate due to the multiple ways of deforming the hexamers, indicating that the experimental realization of the long-range ordered H2O octadecamers might be challenging. Our XRD reveals that the atomic displacement parameters of O are elongated toward the intra-hexamer bond direction (see SM [31]), suggesting that local distortion from the regular hexamer may already occur at 30 K. However, the symmetry lowering from trigonal to orthorhombic structure is not detected within the experimental resolution, probably because 30 K is too close to $T_{\mathrm{c2}}$. Lower-temperature diffraction experiments are needed to discuss the symmetry of the LT’ phase.
Last, we comment on the slight discrepancy of $T_{\mathrm{c1}}$ observed in the dielectric (170–200 K) and XRD (200 K) experiments. The primary order parameter of the HT-LT phase transition is the electric toroidal dipole moment $\mathbf{A}$, which is not a polar multipole. Therefore, the dielectric measurement does not necessarily detect this ordering. The dielectric anomalies observed at $T_{\mathrm{c1}}$ (Fig. 2) probably reflect the ferrotoroidal domain dynamics, since the domain boundaries can locally possess electric dipole moments. Similarly, the deformed hexamers below $T_{\mathrm{c2}}$ possess electric dipole moments, and the alignment of the type-1 hexamers near the domain boundaries can be influenced by external electric fields. The frequency dependences of these dielectric anomalies ($T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$) indicate the energy barrier to move these domain boundaries. Therefore, the thermodynamical transition temperatures are expected to locate at the beginning of the dielectric relaxations, namely, $T_{\mathrm{c1}}\approx 200$ K and $T_{\mathrm{c2}}\approx 50$ K.
In conclusion, we study the water of crystallization on a quasi 2D honeycomb lattice in martyite, which shows the hierarchical ordering at low temperatures. At room temperature (HT), water molecules are dynamically disordered due to the geometrical frustration of the honeycomb lattice. Here, the rotational and vibrational motions are allowed in the porous framework. At $T_{\mathrm{c1}}$, these water molecules cooperatively form H2O hexamers with hydrogen bonds, leading to the ferrotoroidal order in the LT phase. Once the H2O hexamer is formed, the rotational motions of H2O are forbidden, while the vibrational motions along the hydrogen bonds are still alive. Finally at $T_{\mathrm{c2}}$, these vibrational motions of H2O freeze with the deformation of the H2O hexamers in the LT’ phase. After the deformation, the orientation of the type-1 hexamer (location of the long hydrogen bond) appears as a new degree of freedom. The long hydrogen bonds favour lying close to each other, leading to the formation of the H2O octadecamers. These successive phase transitions gradually break the original honeycomb-lattice symmetry and lead the system close to the ground state, which is difficult to achieve on the highly degenerate energy landscape. This degeneracy is the feature of the monolayer ice, behind the polymorphism of H2O.
Acknowledgements
We acknowledge K. Homma, S. Morita, N. Kawashima, D. Hamane, J. Nasu, and S. Hayami for the fruitful discussions. The synchrotron radiation experiments were performed at SPring-8 with the approval of the Japan Synchrotron Radiation Research Institute (JASRI) (Proposal Nos. 2023A1687 and 2024A1709). The crystal structure figures have been created by using the visualization software VESTA [38].
Funding
This work was partly supported by JSPS KAKENHI, Grant-in-Aid for Scientific Research (Nos. 22K14010, 23H04861, 24K06944, and 24H01644).
Author contributions
T.N. conceived and designed the project. T.N., H.I., T.Y., A.M., and Z.H. grew and characterized the single- and polycrystals of martyite. T.N., M.K., O.Y. and T.M. performed the dielectric measurements. T.N., T.H., and N.O. performed the Raman scattering experiments. S.K., Y.N., and T.A. performed the synchrotron XRD experiments on the single crystal. J.K. and K.K. performed the MD simulations. All authors discussed the results. T.N., S.K., and K.K. wrote the manuscript under the supervision of T.H., Z.H., Y.T., T.A., and T.M.
Competing interests
The authors declare no competing interests.
Data and materials availability
The data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
Experimental
Sample preparation. We obtained single- and polycrystals of martyite by hydrothermal synthesis. We put 100 mg of ZnO (99.0+%, Fujifilm Wako) powder, 200 mg of V2O5 (99.0+%, Fujifilm Wako) powder, 200 mg of citric acid, and 50 ml of deionized water in Teflon-lined stainless steel autoclaves, and kept them at 130 C for 10–30 days. Hexagonal transparent single crystals up to the size of 0.5$\times$0.5$\times$0.1 mm3 were obtained.
Dielectric measurement. We measured the dielectric properties of polycrystalline martyite pellets by using an LCR meter (Keysight E4980A). We used silver paint as an electrode and attached the gold wires on both sides of the pellet. To avoid water loss from the crystal, we quickly exchanged the sample chamber with He gas and performed the low-temperature experiment down to 4 K. We note that results were sample dependent, probably because the dielectric properties are sensitive to the water loss and surface adsorbed water. Nevertheless, the two anomalies at $T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$ are reproduced.
Single-crystal XRD. We performed the XRD experiments using a single crystal of 80$\times$70$\times$10 $\mu$m3 on BL02B1 at a synchrotron facility SPring-8 in Japan [39]. We employed a He-gas-blowing device for controlling the temperature between 30 and 300 K. The X-ray wavelength was $\lambda=$ 0.30918 Å. We used a two-dimensional detector CdTe PILATUS to record the diffraction pattern. The intensities of Bragg reflections with the interplane distance $d>0.3$were collected by CrysAlisPro program [40]. Intensities of equivalent reflections were averaged and the structural parameters were refined by using Jana2006 [41].
MD simulation. We carried out molecular dynamics (MD) simulations of two-dimensional water in martyite using GROMACS 2018 software [42]. The model system consists of $N_{w}$ water molecules confined between two atomic layers of martyite. The atomic positions of martyite other than those of water molecules were fixed during the simulations. The number $N_{w}$ of water molecules was 288 unless otherwise stated. The simulation cell was a rectangular prism under periodic boundary conditions in the two dimensions parallel to the atomic layers. The temperature was controlled by the Nosé-Hoover thermostat and the volume of the system was fixed. Other details are described in SM.
Supplemental Material of ”Ferrotoroidal order of the monolayer ice in martyite”
I Crystal growth
We grew single- and polycrystals of martyite by hydrothermal synthesis. We put 100 mg of ZnO (99.0+%, Fujifilm Wako) powder, 200 mg of V2O5 (99.0+%, Fujifilm Wako) powder, 200 mg of citric acid, and 50 ml of deionized water in Teflon-lined stainless steel autoclaves, and kept them at 140 C for 10–30 days. Hexagonal transparent single crystals up to the size of 0.5$\times$0.5$\times$0.1 mm3 were obtained as Fig. S1(a). The edge of the crystal was often red-coloured, probably because of crystallographic defects and water loss. We note that the green part of the photo (Fig. S1(b)) also has a similar crystal structure, although we did not use it for the experiments. The green colour probably indicates that the valence of V ions is shifted from 5+.
Fig. S 1: Photographs of martyite. (a) Single crystals of martyite. (b) Crystal cluster of martyite.
The crystal structure of martyite is topologically the same as the kagomé antiferromagnet, volborthite [Cu3(V2O7)(OH)2$\cdot$2H2O] [43, 44, 45]. The crystal system of volborthite is monoclinic because of the Jahn–Teller distortion of Cu2+ ions. The water of crystal in volborthite is ordered although the molecular orientation changes upon the orbital switching or flipping around 300 K [46]. In contrast, the crystal structure of martyite is trigonal because the $3d$ orbitals of Zn2+ ions are fully occupied and the Jahn–Teller instability is absent. The water of crystal in martyite is located at the perfect honeycomb lattice and is dynamically disordered at room temperature. This difference indicates that the geometrical frustration of the perfect honeycomb lattice suppresses the ordering of H2O molecules.
II Single-crystal XRD
The results of the structural analysis of martyite are summarized in Tables S I–S IV and Figures S2–S5. Figures S2(a) and S2(b) show the crystal structure at 30 K (LT) with the space group $P\overline{3}$. In the structural refinement, we fixed the Ow–Hw bond length and Hw-Ow-Hw bond angle in H2O molecules at 0.95 Å and 104.5°104.5°104.5  °, respectively. Figure S2(c) shows the $|F_{\mathrm{c}}|^{2}$$|F_{\mathrm{o}}|^{2}$ plot resulting from the structural analysis, and Fig. S2(d) displays only the weak superlattice reflections. Here, $F_{\mathrm{c}}$ and $F_{\mathrm{o}}$ indicate the calculated and observed crystal structure factor, respectively. The $|F_{\mathrm{c}}|^{2}$ and $|F_{\mathrm{o}}|^{2}$ of each diffraction reflection show high linearity, indicating good agreement between the experimental and calculated intensities.
Figure S2(e) shows the temperature dependence of the Ow–Ow bond distances in the H2O layer. Figure S2(f) shows the temperature dependence of the superlattice peak intensity. The black dots indicate the observed intensity of -4/3 5/3 0, while the red dots represent the calculated intensity of -4/3 5/3 0, considering only the Ow displacement $\Delta\bm{u}_{\mathrm{O}}$ from the $P\overline{3}m1$ structure. The superlattice peak intensity scales primarily with the displacement of O atoms within the H2O molecule.
If the toroidal arrangement of water molecules is randomly distributed between the nearest hexamers, the structure can be analyzed using a model that assumes disorder of water molecules within the $P\overline{3}1m$ space group. Specifically, this can be represented by a split model for the Ow atoms that constitute the water molecules. We performed structural analysis at 30 K, assuming the $P\overline{3}1m$ space group. The analysis results before and after incorporating the Ow-site split model are shown in Fig. S3(a) and S3(b), respectively. Here, the occupancy of the Ow site is fixed at 1/2. Note that when analyzing a model including hydrogen atoms, the parameters did not converge. Therefore, we present analyses that exclude hydrogen atoms. The split width corresponding to the Ow-Ow distance is 0.161(13) Å. When the Ow-site split model is applied, the $R$ values of the analysis improve; however, it remains significantly worse than that obtained under the assumption of the $P\overline{3}$ space group (Table S II). These results suggest that, in the low-temperature phase, the toroidal order of water molecules is not random but exhibits ferroic arrangement within the plane.
Assuming the $P\overline{3}$ space group, the atomic displacement parameter of Ow in the H2O hexamer at 30 K extends along the bonding direction, as shown in the inset of Fig. S2(b). Therefore, we performed the structural analysis using a site splitting model for Ow. Figure S4 shows the results of the splitting model analysis at 30 K, where the Hw atoms of the H2O molecule have been removed. All structural parameters were converged, with the occupancy ratio of the split oxygen atoms being approximately Ow1:Ow2 = 2:1. Consequently, shorter Ow1–Ow1 and Ow1–Ow2 bond distances were obtained compared to the non-split model, which may be related to the deformation of water hexamers observed in the MD simulations.
Considering the dipole-dipole interactions of water molecules between the layers, an anti-ferroic toroidal arrangement along the interlayer direction is more stable than the ferroic one observed in the experiment (see also Fig. S6(b)). We examine the relationship between the framework structure–composed of ZnO4(OH)2 octahedra and V2O7 pillars–and the interlayer water molecules. Figure S5 presents the crystal structure at 30 K, highlighting the ZnO4(OH)2 layers and V2O7 pillars. The structure contains two Zn sites and three V sites with distinct symmetries. In the V(1)2O7 pillar, the distance between the V atoms above and below the central O is uniform. However, in the V(2)V(3)O7 pillar, the V-O distances above and below the central O differ, indicating an electric dipole moment along the $c$-axis, represented by grey vectors. This dipole moment, originating from the V(2)V(3)O7 pillar, shows a ferroic arrangement along the $c$-axis and appears energetically stable. Notably, the $c$-axis component of the electric dipole moment from the adjacent water molecules opposes that of the V2O7 pillar. This suggests that the ferroic deformation of the V(2)V(3)O7 induces the local dipole moment in the framework and stabilizes the ferroic toroidal arrangement of the water molecules’ dipole moments between the layers.
Fig. S 2: Structural analysis assuming $P\overline{3}$ at 30 K. (a) Crystal structure of martyite at 30 K. (b) Arrangement of H2O molecules between layers. A H2O hexamer within the black dashed line shows the atomic displacement parameters, indicated by ellipsoids. $|F_{\mathrm{c}}|^{2}$$|F_{\mathrm{o}}|^{2}$ plot of the structural analysis at 30 K showing (c) fundamental and superlattice reflections and (d) only superlattice reflections. (e) Temperature dependence of the Ow–Ow bond distances in the H2O layer. (f) Temperature dependence of the superlattice peak intensity of -4/3 5/3 0. The black dots represent the observed intensity, while the red dots represent the calculations that only consider the Ow displacement from the $P\overline{3}m1$ structure in the HT phase.
Fig. S 3: Structural analysis assuming $P\overline{3}1m$ at 30 K. Results of the structural analysis at 30 K, assuming the $P\overline{3}1m$ space group. Figures (a) and (b) show the analysis results before and after incorporating the Ow-site split model, respectively. Atomic positions are represented as spheres, while atomic displacement parameters (ADPs) are depicted as ellipsoids.
Fig. S 4: O site split model analysis assuming $P\overline{3}$ at 30 K. (a) Results of the Ow split model analysis at 30 K. A H2O hexamer within the black dashed line shows the atomic displacement parameters, indicated by ellipsoids.
Fig. S 5: Framework deformation in the ferrotoroidal phase. Crystal structure of martyite at 30 K. The Zn layer changes little during the phase transition, while in the V2O7 pillar, the O atom between the V atoms displaces by up to 1 % along the $c$-axis direction. The electric dipole moment due to the displacement of the O atoms in the V2O7 pillars interacts with the dipole moment of H2O molecules, resulting in a ferrotoroidal order not only in-plane but also in the stacking direction.
Tab. S 1: Summary of crystallographic data of martyite at 300 K (HT).
Wavelength (Å) 0.30918
Crystal dimension ($\mu$m3) 80$\times$70$\times$10
Space group $P\overline{3}m1$
$a$ (Å) 6.0325(3)
$c$ (Å) 7.1667(5)
Chemical formula Zn3V2O7(OH)2$\cdot$2H2O
$Z$ 1
$F(000)$ 230
($\sin\theta$/$\lambda$)max-1) 1.67
$N_{\mathrm{total}}$ 19686
$N_{\mathrm{unique}}$ (all) 3017
$N_{\mathrm{unique}}$ ($I>3\sigma$) 2089
Average redundancy 6.525
Completeness (%) 95.14
$N_{\mathrm{parameters}}$ 25
$R_{1}$ ($I>3\sigma$ / all) (%) 2.78 / 4.82
$wR_{2}$ ($I>3\sigma$ / all) (%) 3.89 / 4.52
GOF ($I>3\sigma$ / all) 1.58 / 1.50
CCDC 2425731
Tab. S 2: Summary of crystallographic data of martyite at 30 K (LT).
Wavelength (Å) 0.30918
Crystal dimension ($\mu$m3) 80$\times$70$\times$10
Space group $P\overline{3}$
$a$ (Å) 10.41250(10)
$c$ (Å) 7.14370(10)
Chemical formula Zn3V2O7(OH)2$\cdot$2H2O
$Z$ 3
$F(000)$ 690
($\sin\theta$/$\lambda$)max-1) 1.67
$N_{\mathrm{total}}$ 117979
$N_{\mathrm{unique}}$ (all) 16135
$N_{\mathrm{unique}}$ ($I>3\sigma$) 10665
Average redundancy 7.312
Completeness (%) 94.38
$N_{\mathrm{parameters}}$ 85
$R_{1}$ ($I>3\sigma$ / all) (%) 3.19 / 5.54
$wR_{2}$ ($I>3\sigma$ / all) (%) 4.23 / 4.69
GOF ($I>3\sigma$ / all) 1.91 / 1.72
Flack parameters ($a,b,c:b,a,c$) 0.4992(18):0.5008
CCDC 2425733
Tab. S 3: Structural parameters of martyite at 300 K (HT). The space group is $P\overline{3}m1$ (No. 164) and $a=6.0325(3)$, $c=7.1667(5)$.
Atom Wyckoff Symmetry $x$ $y$ $z$ Occ. $U_{11}$2) $U_{22}$2) $U_{33}$2) $U_{12}$2) $U_{13}$2) $U_{23}$2)
position
Zn $3e$2/m. 1/2 0 0 1 0.01484(4) 0.00968(3) 0.02778(7) 0.004840(17) -0.002628(16) -0.00526(3)
V $2c$ 3m. 0 0 0.747924(19) 1 0.00935(3) 0.00935(3) 0.00759(3) 0.004673(14) 0 0
O1 $2d$ 3m. 1/3 2/3 0.87724(10) 1 0.01010(11) 0.01010(11) 0.01140(18) 0.00505(5) 0 0
O2 $1b$3m. 0 0 1/2 1 0.0249(4) 0.0249(4) 0.0102(3) 0.01247(19) 0 0
O3 $6i$m. 0.30855(10) 0.15427(5) 0.82608(8) 1 0.01354(13) 0.01379(10) 0.01584(15) 0.00677(7) -0.00234(11) -0.00117(6)
O4w $2d$ 3m. 2/3 1/3 0.4890(3) 1 0.0911(17) 0.0911(17) 0.0169(6) 0.0456(8) 0 0
H1 $2d$ 3m. 1/3 2/3 0.74469(10) 1
H2w $6i$m. 0.5784(15) 0.157(3) 0.521(9) 2/3 (fix)
Tab. S 4: Structural parameters of martyite at 30 K (LT). The space group is $P\overline{3}$ (No. 147) and $a=10.41250(10)$, $c=7.14370(10)$.
Atom Wyckoff Symmetry $x$ $y$ $z$ $U_{11}$2) $U_{22}$2) $U_{33}$2) $U_{12}$2) $U_{13}$2) $U_{23}$2)
position
Zn1 $3e$1 1/2 1/2 0 0.00474(5) 0.00581(6) 0.00876(6) 0.00242(3) 0.00155(4) 0.00175(4)
Zn2 $6g$ 1 0.333132(13) 0.166061(9) 0.005430(10) 0.00457(2) 0.00298(3) 0.005506(17) 0.00152(3) 0.00073(4) -0.00013(2)
V1 $2c$ 3.. 0 0 0.25025(4) 0.00405(4) 0.00405(4) 0.00485(6) 0.002026(18) 0 0
V2 $2d$ 3.. 2/3 1/3 0.25557(2) 0.00377(3) 0.00377(3) 0.00274(3) 0.001885(12) 0 0
V3 $2d$ 3.. 2/3 1/3 0.75317(2) 0.00357(2) 0.00357(2) 0.00220(3) 0.001784(12) 0 0
O1 $1b$3.. 0 0 1/2 0.0110(4) 0.0110(4) 0.0033(4) 0.00552(18) 0 0
O2 $2d$ 3.. 2/3 1/3 0.50240(7) 0.00756(16) 0.00756(16) 0.0050(2) 0.00378(8) 0 0
O3 $6g$ 1 0.51334(7) 0.33407(14) 0.83262(6) 0.00166(13) 0.00889(14) 0.00631(9) 0.00116(14) 0.00027(10) -0.00129(18)
O4 $6g$ 1 0.15510(10) -0.00107(12) 0.17192(12) 0.0108(2) 0.00339(14) 0.0092(2) 0.00068(18) -0.00001(18) -0.00157(17)
O5 $6g$ 1 0.82202(8) 0.33391(14) 0.17835(6) 0.00575(19) 0.00763(13) 0.00676(10) 0.00403(18) 0.00033(12) -0.0019(2)
O6 $6g$ 1 0.33317(7) 0.33491(7) 0.12567(4) 0.00484(10) 0.00441(10) 0.00539(6) 0.00124(14) 0.00163(13) 0.00136(13)
O7w $6g$ 1 0.30981(14) 0.33398(12) 0.49011(9) 0.0809(8) 0.0185(3) 0.00842(14) 0.0243(5) 0.0028(4) 0.0011(3)
H1w $6g$ 1 0.3844(13) 0.4228(12) 0.546(3)
H2w $6g$ 1 0.344(2) 0.2654(17) 0.506(2)
H3 $6g$ 1 0.2942(17) 0.2974(17) 0.2462(11)
III Energy scale of the intermolecular interactions
Fig. S 6: Dipole-dipole interaction energies between water molecules. (a) In-plane nearest-neighbour interactions. (b) Out-of-plane nearest-neighbour interactions. The numbers in the brackets indicate the degeneracy.
The crystal structure of martyite suggests that the water of crystallization is a quasi-2D system where each water layer is separated by the ZnO4(OH)2 block layers. Here, each H2O molecule can be hydrogen-bonded only with the in-plane nearest neighbors. The second important interaction is the electric dipole-dipole interaction. In this section, we evaluate the dipole-dipole interaction energy scale in martyite. For simplicity, we neglect the electric polarization of the framework of martyite and treat the intermolecular medium as vacuum. The intermolecular distances are fixed at the experimental values determined by the single-crystal XRD at 300 K.
Figure S6(a) shows the dipole-dipole interaction energies of the in-plane nearest-neighbour pair, $E_{\mathrm{iNN}}$. Here, we assume that H2O molecules at A and B sites can take only three orientations as suggested by the XRD at 300 K. The out-of-plane displacement of H2O, which is suggested from the low-temperature XRD, is not taken into account. The interaction energy is estimated as $E_{\mathrm{iNN}}=$ $+1170$ K, $+732$ K, $-146$ K, $-585$ K, from top to bottom molecular geometries. Considering the energy scale of hydrogen bonds from $-2000$ K to $-3000$ K, the bottom geometry is further stabilized. Therefore, water molecules tend to increase the number of hydrogen bonds towards low temperatures. When the hydrogen-bonded H2O hexamers are formed, two types of hexamers appear with CW or CCW electric toroidal dipole moment. The inter-hexamer interaction is determined by the molecular geometry between two hexamers. This energy diagram suggests that the second geometry from the bottom ($-146$ K) is much stabler than the third one ($+732$ K). Indeed, the XRD results suggest that ferrotoroidal order is realized in the LT phase, where the inter-hexamer molecular geometry is of this type.
Second, we estimate the dipole-dipole interaction energy of the out-of-plane nearest-neighbour pair, $E_{\mathrm{oNN}}$. Similarly, we assume that H2O molecules can take only three orientations. The interaction energy is estimated as $E_{\mathrm{oNN}}=$ $+67$ K (top) and $-34$ K (bottom). This energy scale is much smaller than the in-plane one, with the typical ratio of 50 K/2500 K = 2 %, including the hydrogen bonds. Therefore, as indicated from the crystal structure, martyite can be considered as a quasi-2D system. We should note, however, that the out-of-plane interactions can be mediated via the 3D framework deformation as Fig. S5. Nevertheless, the agreement between the experiments and the MD simulations on the monolayer ice suggests that the expression ”monolayer ice in martyite” is not just an exaggeration, but captures the features of this mineral.
IV Raman scattering
We obtained Raman scattering spectra on a single crystal of martyite with the size of 0.5$\times$0.4$\times$0.1 mm3 using a quasi-backscattering geometry (see Fig. S7). We used a 532 nm single-frequency laser (Cobolt Samba) with linear polarization for the incident light. We analyzed the scattered light using a triple monochromator (JASCO NR-1800) equipped with a linear polarizer and a CCD detector (Princeton Instruments Inc. LN/CCD-1100PB). Temperature dependence above 80 K was conducted using THMS600 (LINKAM SCIENTIFIC) with liquid-N2 under a vacuum environment, and below 150 K using a cryostat with GM refrigerator under He atmosphere. We measured Raman spectra on the as-grown $c$ surface of martyite with parallel and perpendicular polarization geometries of the scattered light to the incident light.
The upper panel of Fig. S9 shows the Raman scattering spectra of martyite at 295 K and 4 K. The Raman peaks can be assigned as the framework or the H2O molecules’ origins. Based on the numerical simulation, we conclude that the lower-energy peaks from 240 to 900 cm-1 originate from the framework vibration. H2O molecule has three vibrational modes, O-H symmetric streching ($v_{1}$), H-O-H bending ($v_{2}$), and O-H asymmetric streching ($v_{3}$) modes [47]. We assign that the relatively weak peak at 1620 cm-1 is $v_{2}$ mode and the band from 3300 to 3600 cm-1 corresponds to the superposition of $v_{1}$ and $v_{3}$ modes. Although the O-H vibration (the $A_{g}$ representation mode under $\overline{3}m$ point group) of the martyite framework is also Raman active, it was not observed probably due to weak intensity. The asterisks indicate the artefacts due to the grease for sample mounting.
Fig. S 7: Quasi-backscattering geometry of the Raman scattering experiment. The sample is mounted on the aluminium stage with a slit for reducing the scattered light from the aluminium surface. The incident angle is approximately 20°20°20  ° with respect to the $c$ axis of the sample. The polarization of the incident light is parallel to the $ab$ plane of the sample.
Fig. S 8: Raman spectra of martyite. The upper panel shows the Raman scattering spectra at 295 K and 4 K. The schematic vibrational modes of H2O ($v_{1}$, $v_{2}$, and $v_{3}$) are also shown. The asterisks indicate the peaks due to the grease for sample mounting. The lower panel shows the assigned peak energies at each temperature. The red symbols indicate the peaks observed only at lower temperatures.
Fig. S 9: Stretching modes of the Raman spectra. The upper panel shows the enlarged Raman scattering spectra around $v_{1}$ and $v_{3}$ from 285 K to 4 K. The lower panel shows the assigned peak energies (p1–p4) at each temperature. The red and green dashed lines are shown for the guide.
The lower panel of Fig. S9 shows the peak energies at each temperature. Red symbols indicate the peak appearing at lower temperatures. The peaks at 305 cm-1, 774 cm-1, 882 cm-1 appear at around 150 K, where the ferrotoroidal ordering of H2O occurs at $T_{\mathrm{c1}}$ (LT). These new peaks indicate the symmetry breaking of the framework as suggested by the XRD experiment. Among the vibrational modes of H2O, the $v_{2}$ peak does not show any peak shift or splitting. In contrast, $v_{1}$ and $v_{3}$ modes around 3400 cm-1 show complicated peak splitting patterns toward low temperatures.
Figure S9 shows the enlarged Raman spectra of martyite from 285 K to 4 K. The upper panel presents the Raman spectra with red (HT), blue (LT), and black (LT’) curves. HT phase show broad 2 peaks at 3500 cm-1 and at 3390 cm-1, probably corresponding to $v_{1}$ and $v_{3}$. Because of the light reflection at the bottom surface of the transparent crystal, we could not see a clear difference in the parallel and perpendicular polarization of the scattered light. Thus, we could not conclude which peak corresponds to $v_{1}$ and $v_{3}$ modes. Since the energy relationship ($v_{1}<v_{3}$ or $v_{3}<v_{1}$) depends on the potential shape around H2O, it is also challenging to estimate from the numerical calculation. Therefore, we keep this assignment for a future issue.
In the LT phase, the peak at 3500 cm-1 (red dashed line) disappears, and a new peak at 3480 cm-1 (green dashed line) appears. The S/N ratio changed from 125 K to 112 K because we used a different cryostat for the low-temperature experiments. Under the determined $P\overline{3}$ crystal structure, there are 4 Raman active modes for $v_{1}$ and $v_{3}$ vibrations, i.e. $2A_{g}+2E_{g}$. However, the structure contains only one site for H2O molecules. Therefore, the result of the two peaks is reasonable. They are assigned to $v_{1}$ or $v_{3}$ mode of the H2O molecules. The energy difference between $A_{g}$ and $E_{g}$ mode, which is distinguished by the phase factor between adjacent H2O molecules, is small and is obscured by their line widths.
In the LT’ phase (even from higher temperatures around 90 K), new peaks appear at 3310 cm-1 and 3440 cm-1. At least, 4 peaks (3390 cm-1 (p1), 3480 cm-1 (p2), 3310 cm-1 (p3), and 3440 cm-1 (p4)) are observed at 4 K. If we assign the 4 modes to $2A_{g}+2E_{g}$ in the $P\overline{3}$ structure, the new modes must appear as the modes are splitting. However, the p3 is a new peak and does not originate from the p1 peak. Therefore, this result suggests that there are at least two H2O sites, and indicates the symmetry lowering probably due to the deformation of the H2O hexamer below $T_{\mathrm{c2}}$ as suggested from the MD simulation.
V MD simulation
V.1 Simulation conditions
The lateral dimensions of the simulation cell for $N_{w}=288$ are 7.236 nm×$\times$ 6.264 nm. The water model employed in the present study is primarily TIP4P/ICE [33] and, alternatively, TIP4P/2005 [48] is used for comparison. The water-martyite interaction is described by the sum of the pair potentials $\phi_{ij}(r)$ between an interaction site $i$ of the water model and an atom $j$ in martyite:
$\phi_{ij}(r)=4\epsilon_{ij}\left[\left(\frac{\sigma_{ij}}{r}\right)^{12}-\left(\frac{\sigma_{ij}}{r}\right)^{6}\right]+\frac{Z_{i}Z_{j}e^{2}}{4\pi\varepsilon_{0}r},$
where the Lennard-Jones (LJ) size and energy parameters $\sigma_{ij}$ and $\epsilon_{ij}$ are given by the geometric averages. The LJ parameters and charges for the atoms in martyite are listed in Table SLABEL:tb:lj.
Tab. S 5: Lennard-Jones parameters and charges for the atoms of the framework of martyite.
Atom $\epsilon$ [kJ/mol] $\sigma$ [nm] $Z$
H [^1] Those of the OH groups facing the layer of water molecules. 0 0 1
[…]
H 0 0 0
O 0.7749 0.3159 0
V 0.1338 0.2630 0
Zn 0.06241 0.2486 0
Water molecules were first arranged at the honeycomb lattice sites with arbitrary molecular orientations, equilibrated for 1 ns at 5 K with the positional constraints on the water’s oxygen atoms, and then further equilibrated for 10 ns at 300 K without any positional constraints. At 300 K and below, without any positional constraints, the water molecules oscillate and rotate around each lattice point, but do not migrate to other lattice points or interstitials. The temperature $T$ was decreased from 300 K to 1 K in steps of 10 K for $T\geq 10$ K and in steps of 1 K for $T<10$ K. The simulated time at each temperature was 10 ns for $T\geq 10$ K, and 3 ns for $T<10$ K, the first 1 ns being used for equilibration and the rest for data sampling. The MD time step was 1 fs. The cutoff distance for the LJ potentials was 1.0 nm. The Coulomb potentials were treated using the particle mesh Ewald method, with the real-space cutoff distance being the same as the cutoff distance for the LJ potentials. For the cooling process to study structural changes, 5 sets of MD simulations were performed starting from independent initial conditions, and the numerical results are the averages over the 5 sets. To examine the effects of system size on the ferrotoroidal ice structure, one set of MD simulations was performed for the large system ($N_{w}=864$, the lateral dimensions of 10.854 nm $\times$ 12.528 nm, TIP4P/ICE model, $T\leq$ 170 K).
V.2 Movie of simulated water molecules in martyite
The movie of molecular dynamics of water molecules in martyite is available online [http://]. The temperature is lowered from 300 K to 10 K in steps of 10 K. One second in the movie corresponds to 1.2 ps. In the movie, one can observe the transition at $T_{\text{c1}}\approx 190$ K from a state in which water molecules vibrate vigorously and rotate randomly at around honeycomb lattice sites to a state in which they form six-membered rings through hydrogen bonding, followed by the transition at $T_{\text{c2}}$ from a state in which the six-membered rings dynamically change their shapes to a state in which the six-membered rings are fixed in the distorted shapes.
V.3 Comparison of MD simulations with the water models TIP4P/ICE and TIP4P/2005
Fig. S 10: Number of hydrogen bonds per water molecule in martyite as a function of temperature. The water models are TIP4P/2005 and TIP4P/ICE.
Fig. S 11: Potential energies of the systems with ferrotoroidal ice and toroidal ice structures. The initial configuration of the ferrotoroidal ice was artificially constructed, while that of the toroidal ice was that which spontaneously formed during the cooling process of the MD simulation. The error bars are shown at each temperature. The inset shows the results from 300 K to 0 K.
Fig. S 12: Ratio of the deformed hexamers as a function of temperature. The results using (a) TIP4P/ICE and (b) TIP4P/2005 models are shown.
V.3.1 Formation of six-membered rings
Both of the water models give qualitatively the same results for the formation of six-membered rings by hydrogen bonding. Figure S10 shows the number $n_{\text{HB}}$ of hydrogen bonds per water molecule as a function of temperature. In the case of TIP4P/2005, $n_{\text{HB}}$, which is slightly smaller at 300 K, increases most rapidly in the range of 190 to 200 K, which is about 20 K higher than in the case of TIP4P/ICE.
V.3.2 Toroidal and ferrotoroidal structures
When the temperature is lowered from 300 K to low temperatures, e.g. 10 K, ferrotoroidal order does not spontaneously appear. However, the potential energy for the ferrotoroidal structure (prepared by design and then equilibrated at low temperatures) is lower than for the toroidal structure, as shown in Figure S11. This is also the case for TIP4P/2005.
V.3.3 Distributions of the deformed six-membered rings (hexamers)
Figure S12 shows the ratio of deformed hexamers in the ferrotoroidal structure as a function of temperature. In the case of TIP4P/ICE, all the hexamers have the same shape with one long bond at low temperatures. In the case of TIP4P/2005, however, 60% of the hexamers have the shape with two long bonds and 40% have the same shape with one long bond. Because of the coexisting hexamers with one- and two-long bonds, the octadecamer formation is hindered with the TIP4P/2005 model. Therefore, the shapes of the hexamers and their distributions are sensitive to the choice of the two water models.